Calculate debroglie`s wavelength of ball of mass 0.1 Kg moving with velocity 30 metre per second

Grade:11

2 Answers

Adarsh
6 years ago
Dear Imaad,
Total energy of particle = ½.m.v2= 45J   and E=h.c/λ
after putting the value we get ,λ  = 44 x 1028m
Thanks and Regards
Adarsh
Vignesh vijayakumar
6 years ago
De broglie equation is , λ=h/mv(h=plank constant)This is obtained by equating E=mv^2 and E=hf=h/λ(v=velocity,f=frequency)Substituting the values in equation,λ=(6.626×10^-34)÷(0.1×30)=2.21×10^-34So de broglie wave length of ball is appx. 2.21×10^-34